How many grams of pb can be extracted from 380 ml 0.4M pb(no3)3?


1
Expert's answer
2021-01-18T01:53:51-0500


n(Pb(NO3)2) = 0.38×0.4=0.152mol

n(Pb(NO3)2)=n(Pb)=0.152mol

m(Pb)=M×n= 207,2×0.152=31,49g


Answer: m(Pb)=31,49g